OpenStudy (anonymous):

a1=3/2; an+1=(n^2+1)/n(an)

OpenStudy (anonymous):

yeah so if anyone is feeling helpful today i need to finish this e2020 thing ive been doing it all sumnmer and its hard and its straight up ridiculous so help me please!!

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

oh find the first five terms og the given sequence

OpenStudy (anonymous):

of*

OpenStudy (anonymous):

\[a_1=\frac{3}{2}\] \[a_{n+1}=\frac{n^2+1}{na_n}\] like that?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

whoa hold on that was wrong!!

OpenStudy (anonymous):

\[a_2\] here i am thinking that \(n+1=2\) and so \(n=1\) now replace \(n\) by 1 and get \[a_2=\frac{1^2+1}{1\times a_n}\] \[a_2=\frac{2}{1\times \frac{3}{2}}\] \[a_2=\frac{4}{3}\]

OpenStudy (anonymous):

a3=10/3?

OpenStudy (anonymous):

now we repeat the process for \(a_3=a_{2+1}\) that is , replace \(n\) by 2 to get \[a_3=\frac{2^2+1}{2\times a_2}\] \[a_3=\frac{5}{2\times \frac{4}{3}}\] \[a_3=\frac{15}{8}\]

OpenStudy (anonymous):

thats not of the options tho

OpenStudy (anonymous):

hmm maybe i made a mistake let me try again

OpenStudy (anonymous):

okay :D

OpenStudy (anonymous):

well i get the same thing again

OpenStudy (anonymous):

shoot i dont even have a calculator :( i just guess all the time

OpenStudy (anonymous):

for \(n+1=3\) we have \(n=2\) \[2\times \frac{4}{3}=\frac{8}{3}\] also \[2^2+1=5\] and \[\frac{5}{\frac{8}{3}}=\frac{15}{8}\]

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

i have repeated it 3 times starting at the beginning to see if i made an arithmetic error and i do not see one

OpenStudy (anonymous):

its cool ill just move on